# Driving Referring Video Object Segmentation with Vision-Language Pre-trained Models

### v1

@@ Abstract
The crux of Referring Video Object Segmentation (RVOS) lies in modeling dense text-video relations to associate abstract linguistic concepts with dynamic visual contents at pixel-level. Current RVOS methods typically use vision and language models pre-trained independently as backbones. As images and texts are mapped to uncoupled feature spaces, they face the arduous task of learning Vision-Language (VL) relation modeling from scratch. Witnessing the success of Vision-Language Pre-trained (VLP) models, we propose to learn relation modeling for RVOS based on their aligned VL feature space. Nevertheless, transferring VLP models to RVOS is a deceptively challenging task due to the substantial gap between the pre-training task (image/region-level prediction) and the RVOS task (pixel-level prediction in videos). In this work, we introduce a framework named VLP-RVOS to address this transfer challenge. We first propose a temporal-aware prompt-tuning method, which not only adapts pre-trained representations for pixel-level prediction but also empowers the vision encoder to model temporal clues. We further propose to perform multi-stage VL relation modeling while and after feature extraction for comprehensive VL understanding. Besides, we customize a cube-frame attention mechanism for spatial-temporal reasoning. Extensive experiments demonstrate that our method outperforms state-of-the-art algorithms and exhibits strong generalization abilities.
@@
Referring video object segmentation, vision-language pre-trained models, temporal modeling, vision-language relation modeling
@@ I Introduction
Referring Video Object Segmentation (RVOS) aims to segment the target object in a video according to the referring expression. It has a wide range of applications, including language-based robot controlling [1, 2], augmented reality [3], and video editing [4, 5]. As language descriptions inherently exhibit flexibility and diversity, RVOS necessitates comprehensive Vision-Language (VL) understanding abilities to accurately discover and segment the referred target.
@@
The crux of RVOS lies in modeling dense text-video relations to associate the diverse yet abstract linguistic concepts with dynamic visual contents at pixel-level. Massive endeavors [6, 7, 3] have been made for this purpose in the RVOS community. Existing RVOS algorithms [6, 7, 3] typically build relation modeling components on independently pre-trained vision and language backbones, including ResNet [8], Video-Swin [9], and RoBERTa [10]. Such a paradigm for learning relation modeling can be summarized as Figure 1 (a). As the backbones map input images and texts into decoupled feature spaces, these algorithms face the challenge of learning VL relation modeling for RVOS from scratch. Although resorting to sophisticated relation modeling mechanisms, they struggle to understand complicated descriptions and videos.
@@
Recently, Vision-Language Pre-trained (VLP) models, such as CLIP [11] and VLMo [12], which can map images and texts into aligned feature space, have drawn much attention. VLP models have played a pivotal role in advancing various tasks, such as zero-shot classification [11] and referring image segmentation [13, 14]. Nevertheless, the application of VLP models in RVOS remains unexplored. In light of this, we seek to unleash the power of VLP models by transferring them to RVOS, allowing us to learn relation modeling for RVOS based on the aligned VL features instead of from scratch, as shown in Figure 1 (b). Compared with the transfer to image segmentation [13, 15], the transfer to RVOS poses a more formidable challenge due to the substantial gap between the pre-training task (image/region-level prediction) and the RVOS task (pixel-level prediction in videos). Particularly, the transfer to RVOS demands not only adapting the image/region-level representation for pixel-level prediction, but also empowering the pre-trained model with temporal modeling ability. As shown in Figure 2, a model using CLIP [11] without temporal modeling loses the target person when the blue jeans disappear in the $140^{th}$ frame.
In this work, we introduce a framework called VLP-RVOS, with the goal of transferring the knowledge of VLP models to learn robust spatial-temporal and vision-language relation modeling for RVOS. The primary challenge is to learn the task-specific prior knowledge from limited video data without forgetting the pre-trained knowledge of vision-language association. To address the issue, we resort to parameter-efficient prompt-tuning, which keeps the VLP model frozen to retain the pre-trained knowledge, instead of fine-tuning it. In particular, we propose a temporal-aware VL prompt-tuning method, which not only adapts the pre-trained VL features for pixel-level prediction but also empowers the vision encoder to capture temporal clues. For comprehensive VL understanding, our framework incorporates multi-stage VL relation modeling, including 1) introducing the linguistic reference guidance during visual feature extraction, 2) fusing the deep VL features after feature extraction, and 3) VL relation modeling during spatial-temporal reasoning. Besides, we tailor a cube-frame attention mechanism to efficiently and effectively perform spatial-temporal reasoning for RVOS.
Extensive experiments on four benchmarks [16, 17, 18] show that VLP-RVOS outperforms state-of-the-art methods. Figure 3 illustrates the comparison on Ref-DAVIS17 [18] in learnable param and $\mathcal{J}\&\mathcal{F}$. The favorable performance demonstrates that our framework effectively unleashes the power of VLP to RVOS. Our contributions can be concluded as:
  • We introduce the VLP-RVOS framework, facilitating the transfer of VLP models to RVOS; it enables us to learn relation modeling for RVOS from aligned VL feature space instead of from scratch.
  • We propose a temporal-aware prompt-tuning method, which not only adapts pre-trained features for pixel prediction but also enables the vision encoder to capture temporal clues.
  • We propose a multi-stage VL relation modeling scheme for comprehensive VL understanding, and tailor a cube-frame attention mechanism for efficient and effective spatial-temporal reasoning.
@@ II Related Work
Nevertheless, these RVOS algorithms construct the relation modeling components on independently pre-trained vision and language backbones and learn relation modeling from scratch, which is a tough learning task. Unlike these methods, we propose to transfer the powerful VLP model to RVOS, allowing us to learn relation modeling for RVOS from a joint VL feature space instead of from scratch.
Referring image segmentation. Referring Image Segmentation (RIS) is closely related to RVOS, whose goal is to segment the target object described by the referring expression in a static image. Similar to existing RVOS algorithms, numerous RIS approaches [32, 33, 34, 35] adopt a pipeline of first extracting the visual and linguistic features and then modeling the cross-modality relation based on the unimodal representations for image mask prediction. Most of these algorithms [33, 34, 35, 36, 37] resort to independently pre-trained vision and language backbones and learn VL relation modeling from scratch. A few methods [13, 14, 38] build the RIS framework on top of the vision-language pre-trained model, CLIP. Compared with the CLIP-based RIS approaches, transferring VLP models to RVOS is much more challenging due to the larger gap between the pre-training task and the RVOS task.
Vision-language pre-trained models. Recently, VLP models [12, 39, 40, 11, 41], learning multi-modality representation on large-scale image-text pairs, have attracted much attention. They typically adopt a dual-stream [39, 11, 42, 43, 44] or single-stream [40, 41] encoder structure to extract the visual and linguistic features and align them via cross-modality interaction. VLP models have driven the progress of various downstream tasks, such as image-text retrieval [12, 40, 11], referring image segmentation [13, 14, 38], and open-vocabulary detection [45]. Nevertheless, the exploitation of VLP models for RVOS has not been explored. In this paper, we try to overcome the discrepancy between the pre-training and RVOS tasks and take a step towards transferring the powerful VLP models to RVOS.
Prompt-tuning. Prompting was proposed in NLP [46, 47, 48] to generate task-specific instructions for the language model to obtain desired outputs. Recently, prompt-tuning has been widely explored in vision and multi-modal problems to efficiently adapt the pre-trained model to downstream tasks, including image/video recognition [49, 50, 51], image segmentation [14, 52], video-text retrieval [53], and domain adaptation [54]. Nevertheless, prompt-tuning has not been explored in the RVOS area, which requires pixel-level video-text understanding and is different from the above-mentioned tasks. In this work, we explore adapting the pre-trained VL representation to RVOS via prompt-tuning.
@@ III VLP-RVOS
[…]
Figure 4 illustrates the architecture of VLP-RVOS, which mainly consists of the VLP model, the VL Feature Fusion (VLFF) module, and the Spatial-Temporal Reasoning (STR) module. To learn task-specific knowledge from limited video data without forgetting pre-trained knowledge, we opt for the parameter-efficient prompt-tuning method, instead of fine-tuning the VLP model in our framework, which poses the risk of hurting the generalization ability. Specifically, we design a temporal-aware VL prompt-tuning method to enable the vision encoder to capture the temporal clue for video understanding.
For comprehensive VL understanding, VLP-RVOS is designed to conduct three-stage VL relation modeling: 1) We introduce additional Multi-head Cross-Attention (MCA) operations into the vision encoder to exploit the additional linguistic reference guidance for feature extraction. 2) We employ the VLFF module to fuse deep VL features for associating high-level visual semantics with abstract linguistic concepts. 3) We perform VL relation modeling between linguistic features and shallow visual features in the STR module; the goal is to introduce low-level semantic guidance for understanding the motions or appearance variations described in the text. The three stages are marked by the red numbers in Figure 4. Subsequently, we delve into the specifics of VLP-RVOS.
@@ III-B Temporal-aware VL prompt-tuning
To preserve pre-trained knowledge, we opt for prompt-tuning to adapt the VLP model to RVOS, which keeps the VLP model frozen and learns a small number of prompt tokens. Particularly, we design a temporal-aware VL prompt-tuning method to adapt the VLP model for pixel-level prediction and enable it to capture temporal clues. Next, we elaborate on the prompt-tuning method.
@@ III-B1 Temporal-aware vision prompt-tuning
Prompt-tuning on the vision encoder has two objectives: 1) adapting the visual representation pre-trained for image/region-level prediction to pixel prediction; 2) empowering the vision encoder to capture and exploit the temporal clue in videos. To this end, we introduce three types of prompt tokens: the vision prompt, the temporal prompt, and the historical prompt.
Vision prompt. The vision prompt tokens $\bm{P}_{v}\in\mathbb{R}^{M_{v}\times C_{v}}$ are introduced to adapt the pre-trained visual representation for pixel prediction. Technically, they are randomly initialized learnable vectors. We adopt a deep prompt-tuning strategy on the vision decoder to provide additional learning capacity for each transformer layer. Specifically, we divide the $M_{v}$ vision prompt tokens into $L_{v}$ groups, each containing $m_{v}=M_{v}/L_{v}$ prompt tokens. These groups are then appended to the patch tokens of each vision transformer layer. Herein $L_{v}$ is the number of vision transformer layers. All frames share the same prompt tokens in each layer.
@@
Temporal prompt. The temporal prompt tokens $\bm{P}_{tmp}\in\mathbb{R}^{m_{tmp}\times C_{v}}$ are used as carriers to capture and spread the temporal clue in the input video clip. Like $\bm{P}_{v}$, the temporal prompt tokens $\bm{P}_{tmp}$ are also randomly initialized learnable vectors. Differently, we adopt a shallow prompt-tuning strategy for the temporal prompt. We repeat $\bm{P}_{tmp}$ for $T_{c}$ times and append the $t$-th copy $\bm{P}^{t}_{tmp}$ to the patch embeddings of $\bm{I}^{t}$ in the first transformer layer. In each layer, we use the output embeddings corresponding to the temporal prompt tokens as carriers for temporal modeling. Specifically, we construct a Parameter-Reusing Temporal Capture (PRTC) module on the output of each vision transformer layer to capture the temporal clue of the video. In the $l$-th layer, it takes as input the visual features $\mathcal{F}^{l}_{v}\!=\!\{\bm{F}_{v}^{t,l}\}_{t=T}^{T+T_{c}-1}$ and temporal embeddings $\hat{\mathcal{P}}^{l}_{tmp}\!=\!\{\hat{\bm{P}}_{tmp}^{t,l}\}_{t=T}^{T+T_{c}-1}$ of all frames, and outputs the new temporal embeddings $\mathcal{P}^{l}_{tmp}\!=\!\{\bm{P}_{tmp}^{t,l}\}_{t=T}^{T+T_{c}-1}$ modeling the temporal clues. The temporal clues embedded in $\mathcal{P}^{l}_{tmp}$ are further spread to the visual features via the interaction of the $(l\!+\!1)$-th vision transformer layer.
As shown in Figure 5, the PRTC module reuses the frozen visual transformer layer in VLP models as both the encoder and decoder. Technically, we directly replace the MSA operation with the MCA operation to convert an encoder into a decoder. PRTC employs the encoder to perform the interaction between the temporal embeddings of all frames and uses the decoder to perform the interaction between the temporal embeddings and visual features of all frames. The temporal clues are embedded into $\mathcal{P}^{l}_{tmp}$ through the aforementioned cross-frame interactions. Denoting the encoder and decoder by $\Phi^{l}_{Enc}$ and $\Phi^{l}_{Dec}$ in the $l$-th layer, the above operation can be formulated as:
@@
Historical prompt. VLP-RVOS processes long videos clip-by-clip. Therefore, we introduce historical prompt tokens, conditioned on the target states in the previous clip, to provide historical prior for the VLP model. Technically, each historical prompt token is calculated based on the feature of a previous frame through a masked global pooling layer following a linear projection layer. We adopt a deep prompt-tuning strategy with the historical prompts by appending them to every vision transformer layer. Particularly, we use different linear projection layers to generate the historical prompt tokens for each visual transformer layer, as different layers have different semantic levels.
With the above prompt-tuning method, the processing of the $t$-th frame in the $l$-th vision transformer layer can be formulated as:

### v2

# Harnessing Vision-Language Pretrained Models with Temporal-Aware Adaptation for Referring Video Object Segmentation

@@ Abstract
The crux of Referring Video Object Segmentation (RVOS) lies in modeling dense text-video relations to associate abstract linguistic concepts with dynamic visual contents at pixel-level. Current RVOS methods typically use vision and language models pretrained independently as backbones. As images and texts are mapped to uncoupled feature spaces, they face the arduous task of learning Vision-Language (VL) relation modeling from scratch. Witnessing the success of Vision-Language Pretrained (VLP) models, we propose to learn relation modeling for RVOS based on their aligned VL feature space. Nevertheless, transferring VLP models to RVOS is a deceptively challenging task due to the substantial gap between the pretraining task (static image/region-level prediction) and the RVOS task (dynamic pixel-level prediction). To address this transfer challenge, we introduce a framework named VLP-RVOS which harnesses VLP models for RVOS through temporal-aware adaptation. We first propose a temporal-aware prompt-tuning method, which not only adapts pretrained representations for pixel-level prediction but also empowers the vision encoder to model temporal contexts. We further customize a cube-frame attention mechanism for robust spatial-temporal reasoning. Besides, we propose to perform multi-stage VL relation modeling while and after feature extraction for comprehensive VL understanding. Extensive experiments demonstrate that our method performs favorably against state-of-the-art algorithms and exhibits strong generalization abilities.
@@
Referring video object segmentation, vision-language pre-trained models, temporal modeling
@@ I Introduction
Referring Video Object Segmentation (RVOS) aims to segment the target object in a video according to the referring expression. It has a wide range of applications, including language-based robot controlling [1, 2], augmented reality [3], and video editing [4, 5]. As language descriptions inherently exhibit flexibility and diversity, RVOS necessitates comprehensive Vision-Language (VL) understanding abilities to accurately discover and segment the target object.
@@
The crux of RVOS lies in modeling dense text-video relations to associate the diverse yet abstract linguistic concepts with dynamic visual contents at pixel-level. Massive endeavors [6, 7, 3] have been made for this purpose in the RVOS community. Existing RVOS algorithms [6, 7, 3] typically build relation modeling components on independently pretrained vision and language backbones, including ResNet [8], Video-Swin [9], and RoBERTa [10]. Such a paradigm for learning relation modeling can be summarized as Figure 1 (a). As the backbones map input images and texts into decoupled feature spaces, these algorithms face the challenge of learning VL relation modeling for RVOS from scratch. Although incorporating sophisticated relation modeling mechanisms, they struggle to understand complicated descriptions and videos.
@@
Recently, Vision-Language pretrained (VLP) models, such as CLIP [11] and VLMo [12], which map images and texts into aligned feature space, have drawn much attention. They have been pivotal in advancing various tasks, such as zero-shot classification [11] and referring image segmentation [13, 14]. Nevertheless, the application of VLP models in RVOS remains unexplored. In light of this, we seek to unleash the power of VLP models for RVOS, allowing us to learn robust relation modeling for RVOS based on the aligned VL features instead of learning from scratch, as shown in Figure 1 (b). Compared with the transfer to image segmentation [13, 15], the transfer to RVOS poses a more formidable challenge due to the significant gap between the pretraining task (static image/region-level prediction) and the RVOS task (dynamic pixel-level prediction). Particularly, the transfer to RVOS demands not only adapting the image/region-level representation for pixel-level prediction, but also empowering the VLP models with temporal modeling ability. As shown in Figure 2, a model using CLIP [11] without temporal modeling loses the target person when the blue jeans disappear in the $140^{th}$ frame.
@@
In this work, we present a framework called VLP-RVOS which harnesses VLP models for RVOS through temporal-aware adaptation. Specifically, it transfers the knowledge embedded in VLP models to learn robust spatial-temporal and vision-language relation modeling for RVOS. The primary challenge is to learn the above task-specific knowledge from limited video data without forgetting the pretrained knowledge of VL association. To address the issue, we resort to parameter-efficient prompt-tuning, which keeps the VLP model frozen to retain pretrained knowledge and incorporates additional prompts to learn task-specific knowledge. Particularly, we propose a temporal-aware VL prompt-tuning method, which not only adapts the pretrained VL features for pixel-level prediction but also empowers the vision encoder to capture temporal contexts. We also introduce a cube-frame attention mechanism to further facilitate spatial-temporal reasoning for RVOS. Additionally, to ensure comprehensive VL understanding, our framework integrates multi-stage VL relation modeling, including 1) leveraging the linguistic reference to guide visual feature extraction, 2) fusing the deep VL features after feature extraction, and 3) incorporating VL relation modeling during spatial-temporal reasoning.
Extensive experiments on five benchmarks [16, 17, 18, 19] show that VLP-RVOS performs favorably against state-of-the-art methods. Figure 3 illustrates the comparison in learnable param and $\mathcal{J}\&\mathcal{F}$ on Ref-DAVIS17 [18]. Experimental results show that our framework effectively unleashes the power of VLP to RVOS. Our contributions can be concluded as:
  • We present the VLP-RVOS framework harnessing VLP models for RVOS through temporal-aware adaptation. To the best of our knowledge, this is the first framework designed to facilitate robust VL relation modeling for the RVOS task using the VLP models.
  • We propose a temporal-aware prompt-tuning method, which not only adapts pretrained VL features for pixel prediction but also enables the vision encoder to capture temporal contexts.
  • We tailor a cube-frame attention mechanism to facilitate spatial-temporal reasoning for RVOS and propose a multi-stage VL relation modeling scheme for comprehensive VL understanding.
@@ II Related Work
Nevertheless, these RVOS algorithms construct the relation modeling components on independently pretrained vision and language backbones and learn relation modeling from scratch, which is a tough learning task. Unlike these methods, we propose to transfer the powerful VLP model to RVOS, allowing us to learn relation modeling for RVOS from a joint VL feature space instead of learning from scratch.
Referring image segmentation. Referring Image Segmentation (RIS) is closely related to RVOS, whose goal is to segment the target object described by the referring expression in a static image [33, 34]. Similar to existing RVOS algorithms, numerous RIS approaches [35, 36, 37] adopt a pipeline of first extracting the visual and linguistic features and then modeling the cross-modality relation based on the unimodal representations for image mask prediction. Most of these algorithms [36, 37, 38, 39] resort to independently pretrained vision and language backbones and learn VL relation modeling from scratch. A few methods [13, 14, 40] build the RIS framework on top of the vision-language pretrained model, CLIP. Compared with the CLIP-based RIS approaches, transferring VLP models to RVOS is much more challenging due to the larger gap between the pretraining task and the RVOS task.
Vision-language pretrained models. Recently, VLP models [12, 41, 42, 11, 43], learning multi-modality representation on large-scale image-text pairs, have attracted much attention. They typically adopt a dual-stream [41, 11, 44, 45, 46] or single-stream [42, 43] encoder structure to extract the visual and linguistic features and align them via cross-modality interaction. VLP models have driven the progress of various downstream tasks, such as image-text retrieval [12, 42, 11], referring image segmentation [13, 14, 40], and open-vocabulary detection [47]. Nevertheless, the exploitation of VLP models for RVOS has not been explored. In this paper, we try to overcome the discrepancy between the pretraining and RVOS tasks and take a step towards transferring the powerful VLP models to RVOS.
[…]
Prompt-tuning. Prompting was proposed in NLP [48, 49, 50] to generate task-specific instructions for the language model to obtain desired outputs. Recently, prompt-tuning has been widely explored in vision and multi-modal problems to efficiently adapt the pretrained model to downstream tasks, including image/video recognition [51, 52, 53], image segmentation [14, 54], video-text retrieval [55], and domain adaptation [56]. Nevertheless, prompt-tuning has not been explored in the RVOS area, which requires pixel-level video-text understanding and is different from the above-mentioned tasks. In this work, we explore adapting the pretrained VL representation to RVOS via prompt-tuning.
@@ III VLP-RVOS
Figure 4 illustrates the architecture of VLP-RVOS, which mainly consists of the VLP model, the VL Feature Fusion (VLFF) module, and the Spatial-Temporal Reasoning (STR) module. To learn task-specific knowledge from limited video data without forgetting pretrained knowledge, we opt for parameter-efficient prompt-tuning, instead of fine-tuning the VLP model, which poses the risk of hurting the generalization ability. Specifically, we design a temporal-aware VL prompt-tuning method to enable the vision encoder to capture the temporal context for video understanding. Besides temporal-aware prompt-tuning, we also introduce the STR module to enhance the temporal modeling ability for RVOS further.
For comprehensive VL understanding, VLP-RVOS is designed to conduct three-stage VL relation modeling, marked by the red numbers in Figure 4: 1) We introduce additional Multi-head Cross-Attention (MCA) into the vision encoder to leverage the linguistic reference to guide visual feature extraction. 2) We employ the VLFF module to fuse deep VL features for associating high-level visual semantics with abstract linguistic concepts. 3) We perform VL relation modeling between linguistic features and shallow visual features in STR, aiming to introduce low-level semantics to enhance the comprehension of changing visual contents described in the text. Next, we delve into the specifics of VLP-RVOS.
@@ III-B Temporal-aware VL prompt-tuning
To preserve pretrained knowledge, we opt for prompt-tuning to adapt the VLP model to RVOS, which keeps the VLP model frozen and learns a small number of prompt tokens. Particularly, we design a temporal-aware VL prompt-tuning method to adapt the VLP model for pixel-level prediction and enable it to capture temporal clues. Next, we elaborate on the prompt-tuning method.
@@ III-B1 Temporal-aware vision prompt-tuning
Prompt-tuning on the vision encoder has two objectives: 1) adapting the visual representation pretrained for image/region-level prediction to pixel prediction; 2) empowering the vision encoder to capture and exploit the temporal context in videos. To this end, we introduce three types of prompt tokens: the vision prompt, the temporal prompt, and the historical prompt.
Vision prompt. The vision prompt tokens $\bm{P}_{v}\in\mathbb{R}^{M_{v}\times C_{v}}$ are introduced to adapt the pretrained visual representation for pixel prediction. Technically, they are randomly initialized learnable vectors. We adopt a deep prompt-tuning strategy on the vision decoder to provide additional learning capacity for each transformer layer. Specifically, we divide the $M_{v}$ vision prompt tokens into $L_{v}$ groups, each containing $m_{v}=M_{v}/L_{v}$ prompt tokens. These groups are then appended to the patch tokens of each vision transformer layer. Herein $L_{v}$ is the number of vision transformer layers. All frames share the same prompt tokens in each layer.
@@
Temporal prompt. The temporal prompt tokens $\bm{P}_{tmp}\in\mathbb{R}^{m_{tmp}\times C_{v}}$ are used as carriers to capture and spread the temporal context in the input video clip. Like $\bm{P}_{v}$, the temporal prompt tokens $\bm{P}_{tmp}$ are also randomly initialized learnable vectors. Differently, we adopt a shallow prompt-tuning strategy for the temporal prompt. We repeat $\bm{P}_{tmp}$ for $T_{c}$ times and append the $t$-th copy $\bm{P}^{t}_{tmp}$ to the patch embeddings of $\bm{I}^{t}$ in the first transformer layer. In each layer, we use the output embeddings corresponding to the temporal prompt tokens as carriers for temporal modeling. Specifically, we construct a Parameter-Reusing Temporal Capture (PRTC) module on the output of each vision transformer layer to capture the temporal context of the video. In the $l$-th layer, it takes as input the visual features $\mathcal{F}^{l}_{v}\!=\!\{\bm{F}_{v}^{t,l}\}_{t=T}^{T+T_{c}-1}$ and temporal embeddings $\hat{\mathcal{P}}^{l}_{tmp}\!=\!\{\hat{\bm{P}}_{tmp}^{t,l}\}_{t=T}^{T+T_{c}-1}$ of all frames, and outputs the new temporal embeddings $\mathcal{P}^{l}_{tmp}\!=\!\{\bm{P}_{tmp}^{t,l}\}_{t=T}^{T+T_{c}-1}$ modeling the temporal contexts. The temporal contexts embedded in $\mathcal{P}^{l}_{tmp}$ are further spread to the visual features via the interaction of the $(l\!+\!1)$-th vision transformer layer.
As shown in Figure 5, the PRTC module reuses the frozen visual transformer layer in VLP models as its encoder and decoder. Technically, we directly replace the MSA operation with the MCA operation to convert a transformer encoder into a decoder. PRTC employs the encoder to perform the interaction between the temporal embeddings of all frames and uses the decoder to perform the interaction between the temporal embeddings and visual features of all frames. The temporal contexts are embedded into $\mathcal{P}^{l}_{tmp}$ through the aforementioned cross-frame interactions. Denoting the transformer encoder and decoder by $\Phi^{l}_{Enc}$ and $\Phi^{l}_{Dec}$ in the $l$-th layer, the above operation can be formulated as:
@@
Historical prompt. VLP-RVOS processes long videos clip-by-clip. Therefore, we introduce historical prompt tokens, conditioned on the target states in the previous clip, to provide historical prior for the VLP model. Technically, each historical prompt token is calculated by performing masked global pooling and linear projection on the feature of a previous frame with the corresponding mask. We adopt a deep prompt-tuning strategy with the historical prompts by appending them to every vision transformer layer. Particularly, we use different linear projection layers to generate the historical prompt tokens for each visual transformer layer, as different layers have different semantic levels.
With the prompt-tuning method, the processing of the $t$-th frame in the $l$-th vision transformer layer is formulated as: